Ice crystallises in a hexagonal lattice having volume of the unit cell as 132*10-24 cm3.If density of ice is 0.92 g cm-3 at a given temperature , then the number of H2O molecules per unit cell is

Grade:12

2 Answers

Vikas TU
4 years ago
Dear student
We have 
V=132×10^-24 
d= 0.92 
Molecular mass of water =18
Note that here molecular mass is of water molecule 
Using d=z×m÷V×6.02×10^23
Ans is 4 molecule 
Here z is molecules of water because we take m as molecular mass of water.
Good Luck 
TEJA KRISHNA
4 years ago
 
Dear student
We have 
V=132×10^-24 
d= 0.92 
Molecular mass of water =18
Note that here molecular mass is of water molecule 
Using d=z×m÷V×6.02×10^23
Ans is 4 molecule 
Here z is molecules of water because we take m as molecular mass of water.
Good Luck 
****************************************************************ALL THE BEST************************************************************************